OpenStudy (anonymous):

what is 0.26 written decimal as a fraction in simplest form

OpenStudy (anonymous):

13/50

OpenStudy (triciaal):

to convert from decimal to fraction put the 1 under the point and add a zero for each digit 0.26 = 26/100 divide by 2 to reduce answer was given above.

OpenStudy (mathstudent55):

First, read the number. Since there are two decimal places to the right of the decimal point, this number is in the hundredth place. The number is read as twenty-six hundredths. Now write the number as a fraction and reduce.

OpenStudy (mathstudent55):

\(\dfrac{26}{100} = \dfrac{2 \times 13}{2 \times 50} = \dfrac{13}{50}\)